Hello Again:
Thanks for your help. I appreciate it. Here is another equation that I need help on. I am trying hard to get to the root of the answer. Please help me.
Solve the systems of linear equation:
5x - 2y = 1
2x = y + -4
Please show me how to do this. I keep getting stuck. I tried the substitution method. Thank you again.

Answer by pwac(253)

5x-2y=1----1
2x=y-4-----2 (note:y+ -4 can be written y-4)
Re-arrange equ.2 to get
y=2x+4-----3 Substitute y=2x+4 into equ.1
5x-2(2x+4)=1 Expand bracket by multipying by -2
5x-4x-8=1
x-8=1 Subtract-8 from both sides to get x
x=1-(-8)
x=9 Put x=9 into equ.3 to find y
y=2X9 +4 =18+4
y=22
To check use values of x and y in either equation 1 or 2.Let's use 1
5X9 - 2X22=1
45-44 does indeed =1 Hope this helps
Pete